OpenStudy (anonymous):

4.Given that y=2x^2+3x,find the range of values of x for which y>9.pls check my answer and working whether i'm right. @ganeshie8

OpenStudy (anonymous):

\[9=2x^2+3x\]\[9<2x^2+3x\]\[0<2x^2+3x-9\]\[2x^2+3x-9>0\]\[(x+3)(2x-3)>0\]

OpenStudy (kamille):

sorry, how 1.5 can be smaller than -3?

OpenStudy (kamille):

it should be written like -3<x<1.5 so the answer is \[x \in (-3;1.5)\]

OpenStudy (jhannybean):

This inequality states "x is less than -3 but greater than 3/2" It should be saying "x is less that 3/2 but greater than -3" How would you write that? @MARC_

OpenStudy (anonymous):

can u show me the working from the start? @Jhannybean

OpenStudy (jhannybean):

\[\begin{align} 2x^2 +3x >9~\text{ Since} ~~y >9 \\&: ~ 2x^2 +3x -9>0 \\&: ~ x^2+3x-18 > 0 \\&: ~ ( x+6)(x-3)>0 \\&: ~ \left(x + \frac{6}{2}\right)\left(x-\frac{3}{2}\right)>0 \\&: ~ (x+3)(2x-3) >0 \\&: ~ x+3 > 0 \implies x > -3 \\&: 2x-3 > 0 \implies x > \frac{3}{2} \\&: ~ \boxed{ -3 > x >\dfrac{3}{2}} \end{align}\]

OpenStudy (jhannybean):

BUM method: \[\color{red}{2}x^2 +3x \color{red}{-9}\]\[x^2 +3x -8\] Multiply the red colored numbers togeher to simplify the quadratic. That's how I solve these types of quadratics, it's not necessary.

OpenStudy (anonymous):

BUM is what?

OpenStudy (jhannybean):

Whoop,I made a typo, that is not \(-8\) it's supposed to say \(-18\) @MARC_

OpenStudy (jhannybean):

Im not sure exactly :P

OpenStudy (jhannybean):

Any more questions? @MARC_

OpenStudy (anonymous):

Nope...thnx @Jhannybean @Kamille

OpenStudy (jhannybean):

Awesome :D

OpenStudy (anonymous):

@ganeshie8 do u know what is BUM?

ganeshie8 (ganeshie8):

\[(x+3)(2x-3)>0 \] your factorization is good so far, next all you need to is think of a parabola

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

what are the x intercepts of the corresponding parabola ?

OpenStudy (anonymous):

-3 and 3/2

ganeshie8 (ganeshie8):

Yes! and your inequality is > 0 so :